my paper tells me to solve the system
-3y+x=6
6Y-5x=12
i tried and i got (0,2)
but i got stuck and i dont know for sure.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

solve the system
-3y+x=6
6Y-5x=12
-------------
Multiply thru the 1st Eq:
-6y + 2x = 12
6y - 5x = 12
-----------------
Add the two Equations and solve for "x"
-3x = 24
x = -8
-------------------------
Substitute to solve for "y":
-3y -8 = 6
-3y = 14
y = -14/3
==========================
Cheers,
Stan H.

Answer by solver91311(24713)



If (0,2) is an element of the solution set of the system:



Then substituting 0 for and 2 for in either equation should result in a true statement. So let's try it out.





So (0,2) is NOT an element of the solution set of the first equation, hence it is not an element of the solution set of the system. Back to the drawing board.

Multiply the first equation by 5



Now add the two equations:





Go back to the original two equations and multiply the first one by 2:



Add the two equations:





Hence the solution set consists of the single element:

Check:



: True



: True

Both check, so answer checks.


John